Find the value of x so that the area under the normal curve between µ and x is 0.4525 and the value of x is less than µ.

1
Expert's answer
2021-12-07T09:07:59-0500

Let x be a continuous random variable that follows a normal distribution with a mean of and a standard deviation of 25. Find the value of x so that the area under the normal curve between µ and x is 0.4525 and the value of x is less than µ.

"P(x< X < \mu) = 0.4525 \\

P(x < X < 200) = 0.4525 \\

P(X<200) -P(X<x) = 0.4525 \\

P(Z< \frac{200-200}{25}) -P(Z< \frac{x-200}{25}) = 0.4525 \\

P(Z< 0) -P(Z< \frac{x-200}{25}) = 0.4525 \\

0.5 -0.4525 = P(Z< \frac{x-200}{25}) \\

P(Z< \frac{x-200}{25}) = 0.0475 \\
\frac{x-200}{25} = -1.6696 \\

x -200 = 25 \times (-1.6696) \\
x = 200 - 41.74 \\
x = 158.26"